Yesterday,the price of envelopes was 3$ a box, and julie is now willing to buy 8 boxes. Is julie's demand for envelopes elastic or inelastic ? What is julie's elasticity of demand ?
1
Expert's answer
2020-12-21T10:58:14-0500

we can compare prices and quantity for yesterday and today.

Assume yesterday, price at "\$3," quantity demanded = 10 boxes

Again, assume today price increased to "\$3.75," quantity demanded =8 boxes

To find Julie's elasticity of demand, we need to divide the percent change in quantity by the percent change in price.

"\%" change in quantity "=(8-10)/10\times100\%=-20\%"

"\%" change in price "=(3.75-3)/3\times 100 \%=25\%"

"Elasticity =|-20\%/25\%|=|-0.8|= 0.8"

Her elasticity of demand is the absolute value of -0.8, or 0.8. Julie's elasticity of demand is inelastic, since it is less than 1.